Problem 16-09 a-b Shamrock Corporation is preparing the comparative financial statements to be included in the annual report to stockholders. Shamrock employs a fiscal year ending May 31. Income from operations before income taxes for Shamrock was $1,277,000 and $724,000, respectively, for fiscal years ended May 31, 2021 and 2020. Shamrock experienced a loss from discontinued operations of $421,000 on March 3, 2021. A 20% combined income tax rate pertains to any and all of Shamrock Corporation’s profits, gains, and losses. Shamrock’s capital structure consists of preferred stock and common stock. The company has not issued any convertible securities or warrants and there are no outstanding stock options. Shamrock issued 41,000 shares of $100 par value, 6% cumulative preferred stock in 2017. All of this stock is outstanding, and no preferred dividends are in arrears. There were 951,600 shares of $1 par common stock outstanding on June 1, 2019. On September 1, 2019, Shamrock sold an additional 402,000 shares of the common stock at $16 per share. Shamrock distributed a 20% stock dividend on the common shares outstanding on December 1, 2020. These were the only common stock transactions during the past 2 fiscal years. Determine the weighted-average number of common shares that would be used in computing earnings per share on the current comparative income statement for:

Weighted-average number of common shares (1) The year ended May 31, 2020. (2) The year ended May 31, 2021. LINK TO TEXT

Starting with income from operations before income taxes, prepare a comparative income statement for the years ended May 31, 2021 and 2020. The statement will be part of Shamrock Corporation’s annual report to stockholders and should include appropriate earnings per share presentation. (Round earnings per share to 2 decimal places, e.g. $2.55.)
Answer #1

Calculation of Weighted Number of Shares-

Note : If we have to calculate EPS then only we can use Common share as weighted number of shares and the value it has been issued in no way will impact calculation of weighted number of shares

Answer :

Weighted Average Number of common share , Pls refer the below table

Date Shares Outstanding No. of Months (a) outstanding from 31st May 20 Total Month (b) Weights(a/b) Weighted Avg as on 31st May 20 (Shares Outstanding* Weight)
1.6.2019 951600 12 12 1            951,600
1.9.2019 402000 9 12 0.75            301,500
Weighted Avg Number of Common Share as on 31s May 20        1,253,100
Date Shares Outstanding No. of Months (a) outstanding from 31st May 21 Total Month (b) Weights(a/b) Weighted Avg as on 31st May 21 (Shares Outstanding* Weight)
1.Jun.20 1253100 12 12 1        1,253,100
1.Dec.20 250620 6 12 0.5            125,310
Weighted Avg Number of Common Share as on 31s May 21        1,378,410

Pls note the calculation for Shares as of 1.Dec.2020 , It is calculated as 20% of 1253100 as it was the share outstanding on 1.Dec.2020

Answer 2

Pls find below EPS calculated and refer the table and the note

Comparative Income Statement
As on 31st May'21 ($) As on 31st May'20 ($)
Income Before taxes (a)                   1,277,000                      724,000
Loss from discontinued Operations (b)                      421,000
Net Income before taxes (c)= (a)-(b)                      856,000                      724,000
Taxes @20% (d)= ©*20%                      171,200                      144,800
Distriutable Profit ( e )= c-d                      684,800                      579,200
6% preference shares (Note 1) (f)                      246,000                      246,000
Distributable profit for Commons Stock g= (e-f)                      438,800                      333,200
Weighted Number of shares calculated in above answer                   1,378,410                   1,253,100
EPS = (g/No. of shares)                            0.32                            0.27
Note 1
Dividend of preference share
Particulars 31st May 21 31st May 20
No. of shares( a)                        41,000                        41,000
Value of each shares ($) (b )                             100                             100
Total Value c= (a*b)                   4,100,000                   4,100,000
Preference dividend @ 6% of c                      246,000                      246,000